Question about ACID/BASE. Equal volumes of 0.230 M weak base (Kb = 4.0× 10–9) and 0.230 M HCl are mixed. Calculate the pH of the resulting solution. Please show all your work.
Answer #1

Assume weak base is BOH

BOH + HCl ----> BCl + H2O

If volume is equal then concentration becomes half. C= 0.115 M

It undergoes salt hydrolysis

It is salt of weak base+strong acid

pKb = -log(Kb) = 8.40

Now pH

pH = 1/2(14-pKb - log[C])

= 1/2(14-8.4-log(0.115))

= 3.3

pH = 3.3